Find y as a function of t

 if 16y′′ + 136y′ +289y = 0,

y(0) = 7 , y′(0) = 3.



1
Expert's answer
2021-07-19T05:46:34-0400
"16y'' + 136y' +289y = 0"

The characteristic equation


"16r^2+136r+289=0"

"(4r+17)^2=0"

"r_{1,2}=-\dfrac{17}{4}"

The general solution of the homogeneous equation is


"y(t)=c_1e^{(-17\/4)t}+c_2te^{(-17\/4)t}"

"y(0)=7:c_1e^{(-17\/4)(0)}+c_2(0)e^{(-17\/4)(0)}=7"

"c_1=7"

"y=7e^{(-17\/4)t}+c_2te^{(-17\/4)t}"

"y'=-\dfrac{119}{4}e^{(-17/4)t}+c_2e^{(-17/4)t}-\dfrac{17}{4}c_2te^{(-17/4)t}"

"y'(0)=3:-\dfrac{119}{4}+c_2-0=3"

"c_2=\dfrac{131}{4}"

The solution of the given Initial Value Problem is


"y(t)=7e^{(-17/4)t}+\dfrac{131}{4}te^{(-17/4)t}"